OpenStudy (anonymous):

What is the center and radius of (x - 6)2 + (y + 1)2 = 81

OpenStudy (mayankdevnani):

\[\huge Welcome~To~Open~Study\] @2413

OpenStudy (luigi0210):

the radius is whatever the equation equals, just squared so: \[\sqrt{81}=?\] And yes, welcome

OpenStudy (luigi0210):

how do you make that big text? >.<

OpenStudy (whpalmer4):

@Luigi0210 select the text (or fancy math expression), then right click and select Show Math As >TeX commands...

OpenStudy (whpalmer4):

@2413 the formula for a circle with radius \(r\) and center \((h,k)\) is \[(x-h)^2+(y-k)^2=r^2\]Does that look similar to anything you have?

OpenStudy (anonymous):

I'm using iPad

OpenStudy (whpalmer4):

very painful doing OpenStudy on an iPad, I know!

OpenStudy (jhannybean):

The center is the vertex : (h,k) \[\large (x-h)^2 + (y-k)^2 = r^2\]your equation \[\large (x -\color{red}{6})^2+(y-\color{red}{(-1)})^2 = 81\]

OpenStudy (luigi0210):

and the radius is there at the end as r^2

OpenStudy (whpalmer4):

and you can think of that \(r^2\) as \((\sqrt{81})^2\) so \(r = \sqrt{81}\)

OpenStudy (luigi0210):

I actually feel like Luigi right now >.<

OpenStudy (whpalmer4):

what do you feel like when you aren't feeling like Luigi?

OpenStudy (luigi0210):

A normal person that isn't being ignored

OpenStudy (whpalmer4):

Looks like we're all being ignored

OpenStudy (luigi0210):

I summon the powers of negative zone! time to dance :D